OpenStudy (anonymous):

I need help on this question: A rain cloud contains 2.66 X 10^7 kg of water vapor. How long would it take for a 2.00kW pump to raise the same amount of water to the cloud's altitude, 2.00 km?

OpenStudy (anonymous):

m = 2.66 X 10^7 and P = 2.00 kW and h = 2.00 km = 2000 m Since Work = PE and PE = mgh we can use the formula P = W/Dt 2000 W = ((2.66 X 10^7)(9.8)(2000)/t) t = 260680000 = 2.61 X 108 s